OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). A) y = four thirds x + 8 B) y = four thirds x + 3 C) y = −three fourths x + 8 D) y = −three fourths x + 3

OpenStudy (anonymous):

HELP

OpenStudy (luigi0210):

Get the equation in slope-intercept form/

OpenStudy (anonymous):

y=4+4x-3?

OpenStudy (anonymous):

ohh is it y=4+4x/3??

OpenStudy (anonymous):

4x - 3y = -24

OpenStudy (anonymous):

https://www.desmos.com/calculator